OpenStudy (anonymous):

What is the distance between the points (-2,1) and (3,7)?

OpenStudy (anonymous):

\[A. \sqrt{11} B.\sqrt{37} C.\sqrt{61} D.\sqrt{89}\]

OpenStudy (diyadiya):

\[\sqrt{(x2-x1)^2+(y2-y1)^2}\]

OpenStudy (anonymous):

do u know if it is a,b,c, or d

OpenStudy (diyadiya):

x1=-2 x2=3 y1=1 y2=7 can you do now ?

OpenStudy (anonymous):

yeah thanks (:

OpenStudy (diyadiya):

anytime ^_^